OpenStudy (anonymous):

What is the slope of a line parallel to 10x - 5y = 8?

OpenStudy (anonymous):

get y by its self first

OpenStudy (anonymous):

Convert it to slope-intercept form.

OpenStudy (anonymous):

10x - 5y = 8 Subtract 10x to both sides, can you do that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so then it will be -5x=-10x+8

OpenStudy (anonymous):

10x-5y=8 -10x -10x -5y=-10x+8 divide by -5 y= -10x + -8/5

OpenStudy (anonymous):

then x=2x+8 right?

OpenStudy (anonymous):

did the question ask for X?

OpenStudy (anonymous):

Correct! But we have: y = 2x + 8/-5

OpenStudy (anonymous):

because y intercept form y=mx+b

OpenStudy (anonymous):

You accidentally put 'x' in the beginning.

OpenStudy (anonymous):

So now it's in slope-intercept form. y = mx + b Where 'm' is the slope. So what's the slope here? @kylefoster58

OpenStudy (anonymous):

y do we have a 2x?

OpenStudy (anonymous):

We divided -5 to both sides.

OpenStudy (anonymous):

cause 5 goes into 10 2 times

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

-5y = -10x + 8 That's every term on both sides. -5y/-5 = -10/-5x + 8/-5 y = 2x + -8/5

OpenStudy (anonymous):

Yep, so what's the slope of y = 2x - 8/5?

OpenStudy (anonymous):

so the answer is -5 right?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

No

OpenStudy (anonymous):

is it 2?

OpenStudy (anonymous):

what does m stamd for

OpenStudy (anonymous):

Yep

OpenStudy (anonymous):

2 is our slope..and parallel lines have the same slope, so that's ou ranswer.

OpenStudy (anonymous):

*our answer

OpenStudy (anonymous):

just positive 2 right?

OpenStudy (anonymous):

just remember always get y by its self

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\sf y = \color{red}mx + b\) \(\sf y = \color{red}2x - \dfrac{8}{5}\)

OpenStudy (anonymous):

Yep, positive 2.

OpenStudy (anonymous):

What is the slope of a line parallel to 4x + y = -1?